# Forgetting is Everywhere

### v1

@@ Abstract
A fundamental challenge in developing general learning algorithms is their tendency to forget past knowledge when adapting to new data. Addressing this problem requires a principled understanding of forgetting; yet, despite decades of study, no unified definition has emerged that provides insights into the underlying dynamics of learning. We propose an algorithm- and task-agnostic theory that characterises forgetting as a lack of self-consistency in a learner’s predictive distribution over future experiences, manifesting as a loss of predictive information. Our theory naturally yields a general measure of an algorithm’s propensity to forget. To validate the theory, we design a comprehensive set of experiments that span classification, regression, generative modelling, and reinforcement learning. We empirically demonstrate how forgetting is present across all learning settings and plays a significant role in determining learning efficiency. Together, these results establish a principled understanding of forgetting and lay the foundation for analysing and improving the information retention capabilities of general learning algorithms. Project page: https://ben-sanati.github.io/forgetting-is-everywhere-project/.
@@ 1 Introduction
Forgetting is a ubiquitous yet poorly understood phenomenon in machine learning (catainf). When a learner updates its beliefs based on new observations, it often forgets prior knowledge, leading to a degradation in performance on previous observations. Although this behaviour is well documented in continual learning (CL; overcomeCF, overcomeCF) and reinforcement learning (RL; atkinson2021pseudo, atkinson2021pseudo; khetarpal2022towards, khetarpal2022towards), it is not confined to non-stationary settings. Even under standard training regimes with independent and identically distributed (i.i.d.) data, models can forget previously acquired abilities, reducing learning efficiency and stability (chunking).
Classical learning theory offers limited tools to analyse forgetting. Concepts such as bias, variance, and generalisation describe performance at fixed times, typically used at convergence. Forgetting, by contrast, is inherently temporal; it concerns what a model retains, what it discards, and how it balances adaptation with stability over time (jagielski2022measuring).
Most existing studies of forgetting come from the CL literature, where metrics typically track performance degradation on earlier tasks after training on later tasks (chaudhry2018riemannian). Although these measures are widely adopted, they poorly capture forgetting and frequently conflate two distinct phenomena: backward transfer, where new learning improves performance on past tasks (benavides2022theory), and forgetting, where updates erode prior knowledge (jagielski2022measuring), making it challenging to distinguish between constructive and destructive adaptation.
In this paper, we introduce a new conceptual foundation for understanding forgetting, built on the following insight:
If a learner updates its predictions on data it already expects, such an update cannot represent the acquisition of new information; it must instead represent the loss of previously acquired knowledge.
This allows us to characterise forgetting in a precise and general manner. We ask: What is forgetting? When and why does it occur? How does forgetting impact learning?
@@
    Formulation of learning over time: Inspired by hutter2005 and dong2022simple, we define a general theoretical formulation for reasoning about how learners acquire, retain, and lose capabilities during learning (§ 3).
@@
    Conceptualisation of forgetting: Building on our insight, we formulate forgetting from a predictive perspective: as a property of how a learner’s predictions evolve over time. Specifically, we define forgetting as a violation of self-consistency in the learner’s predictive distribution, providing a unified probabilistic foundation that generalises prior conceptions of forgetting (§ 4.2).
@@
    Operational measure: This conceptualisation naturally yields a measure of the propensity to forget, which we propose to assess the validity of our formalism (Definition 4.2).
@@
    Empirical characterisation: We empirically investigate the propensity to forget across diverse environments and learning paradigms, including regression, classification, generative modelling, CL, and RL. The results confirm that the observed forgetting dynamics conform to the expected theoretical characteristics of forgetting and further reveal a trade-off between forgetting and training efficiency, with deep RL being particularly susceptible to forgetting (§ 5).
@@ 2 Related work
Forgetting in supervised settings.
Forgetting is often studied in the context of CL (de2021continual; wang2024comprehensive). In this literature, various proxy metrics are used to measure forgetting in settings where a learner encounters data sequentially from different distributions. Here, forgetting is calculated as the change in performance on previously observed distributions (overcomeCF; chaudhry2018efficient; chaudhry2018riemannian).
However, this approach cannot distinguish between two distinct effects: backwards transfer, where learning on new data improves performance on past tasks (benavides2022theory), and forgetting, where learning on new data degrades previously acquired knowledge (catainf; jagielski2022measuring). Performance on previous tasks depends on both how current data informs the learner about past data (backwards transfer) and the extent to which updating on new tasks results in a loss of previous knowledge. Furthermore, most metrics are limited to CL settings and do not generalise to all learners or domains, motivating the need for a more general definition of forgetting.
Various approaches have sought to quantify forgetting. In toy settings with known data distributions, forgetting can be measured relative to an oracle (lee2021continual). Others examine changes in internal representations (kim2025understanding) or model the trade-off between generalisation and forgetting (raghavan2021formalizing). Meanwhile, mechanistic accounts in associative memory models demonstrate how interference between correlated patterns leads to retrieval errors (hopfield_discrete; Amit). A related line of work studies models trained on their own generated outputs, typically in the context of generative model collapse or stability (alemohammad2023self; shumailov2023curse; bertrand2023stability; scholten2025model). These studies implicitly touch on the phenomenon we formalise as forgetting: updating a model on self-generated targets can degrade previously encoded knowledge or induce systematic shifts in its predictions.
@@
Forgetting has also been studied in continual RL; however, it has received less systematic study than in supervised CL. Early discussions of continual RL already noted the risk of knowledge degradation when agents adapt over long time horizons (ring1994continual; ring1997child). More recent surveys highlight that catastrophic forgetting is a persistent challenge in RL, where continual adaptation to changing state distributions or reward signals can interfere with previously acquired knowledge (khetarpal2022towards; kumar2025continual).
Empirical studies document several forms of forgetting. Value-based methods with function approximation often lose performance on earlier estimates (mnih2015human; van2018deep). This is closely related to the phenomenon of policy churn identified by (schaul2022phenomenon) in which the greedy policy of a value-based learner changes in a large portion of the input space after just a few updates. Policy gradient methods are also prone to overwriting earlier strategies during continual adaptation (ring1994continual; kirkpatrick2017overcoming). Replay buffers are widely used to mitigate these effects by reintroducing past experiences during training (random; fedus2020revisiting). However, such methods only partially alleviate forgetting without addressing its root cause, and thus are insufficient solutions in settings like constrained continual RL (dong2022simple; kumar2025continual).
Unlike in supervised CL, where task boundaries are externally defined and forgetting is measured via performance drops on prior tasks, in RL, the data distribution evolves continuously with the learner’s own policy. As a result, forgetting may manifest as interference with previously learned predictions and the gradual disappearance of coverage over regions of the state space that the agent no longer visits (sutton1998reinforcement). Despite these differences, metrics for measuring forgetting in RL remain underdeveloped, and most approaches borrow from supervised CL.
In this paper, we propose a conceptualisation of forgetting that is broadly applicable across learning paradigms and provides a principled distinction between forgetting and related phenomena, such as backwards transfer. Making this distinction is crucial for understanding whether learner updates reflect knowledge loss or beneficial reuse of information. We further demonstrate the generality of our conceptualisation and show how it supports an empirical measure of forgetting that aligns with theoretical expectations.
@@ 3 Learning and inference processes
We present a general framework in which supervised learning, RL, and generative modelling are all specific cases of a single stochastic interaction process. The formalism draws inspiration from the agent-environment perspective from general RL (hutter2005; Lattimore2014; dong2022simple; Abel2023; hutter2024; kumar2025continual), but is adapted to emphasise conventions standard in machine learning (bishop2006pattern; goodfellow2016deep).
@@ 3.1 Preliminaries: Learning interaction
To formalise this interaction, we introduce the notion of an environment-learner interface, which provides a structure for the data exchanged between them at each time step.
@@
Elements of $\mathcal{X}$ are observations, encompassing the signals the environment emits (e.g., features, targets, rewards, etc.), and elements of $\mathcal{Y}$ are outputs, encompassing the signals emitted by the learner (e.g., actions, classifications, regression values, etc.).
@@
We refer to an element of $\mathcal{H}$ as a history $H\in\mathcal{H}$, which is a sequence $H_{0:t}=((X_{0},Y_{0}),\dots,(X_{t},Y_{t}))$ of $t+1$ observation-output pairs $(X_{i},Y_{i})$. For a sequence $A=(A_{0},A_{1},\dots,A_{t})$ and indices $0\leq i\leq j\leq t$, we denote the subsequence between indices $i$ and $j$ as $A_{i:j}=(A_{i},A_{i+1},\dots,A_{j})$.
@@
A learner relative to interface $(\mathcal{X},\mathcal{Y})$ is a tuple $(\mathcal{Z},f,u,u^{\prime},p_{Z_{0}})$, consisting of:
@@
Although a learner’s entire state could, in principle, be updated by a single function (as in work by dong2022simple, dong2022simple and kumar2025continual, kumar2025continual), we distinguish between two update functions to capture different modes of learner evolution. During interaction with the environment, the learning-mode update $u$ governs the evolution of the entire state, including predictive parameters and auxiliary components, such as replay buffers. In contrast, the inference-mode update $u^{\prime}$ allows auxiliary components of the state, such as buffers or counters, to evolve while keeping the predictive parameters fixed. This distinction allows an observer to analyse the learner’s behaviour in both dynamic conditions, where beliefs are updated over time, and in static conditions, where beliefs remain fixed.
@@
The (learning-mode) interaction process between an environment $(e,p_{X_{0}})$ and a learner $(\mathcal{Z},f,u,u^{\prime},p_{Z_{0}})$ relative to an interface $(\mathcal{X},\mathcal{Y})$ is the stochastic process over $(X_{t},Y_{t},Z_{t})$ defined by:
Initialisation: $Z_{0}\sim p_{Z_{0}},\quad X_{0}\sim p_{X_{0}},$
Interaction: $Y_{t}\sim q_{f}(\cdot\mid Z_{t-1},X_{t-1}),$ (learner samples output)
$X_{t}\sim p_{e}(\cdot\mid H_{0:t-1},Y_{t}),$ (environment samples current observation)
@@
This interaction process generates the stochastic processes $X_{t}\,(t\geq 0),\,Y_{t}\,(t>0),\,\text{and }Z_{t}\,(t\geq 0)$.
3.2 Induced future distributions
Figure 1: State evolution and induced futures. The learner’s internal state $Z_{t}$ evolves through training updates $u$ given $(Z_{t-1},X_{t},Y_{t})$. Each state induces a distribution over induced futures $q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t})$, which is updated in inference mode $u^{\prime}$. This separation illustrates how each state encodes retained capabilities, and how training vs. introspective updates interact.
At any time during learning, the learner’s state encodes expectations about how future interactions will unfold. We can make these expectations explicit by asking the learner to internally simulate the future. Rather than waiting for new observations, we let the learner “roll forward” in inference mode using update $u^{\prime}$ to project how the sequence of inputs and outputs may continue. This is very similar to placing a learner into inference mode, where it projects how the sequence of outputs might continue.
We refer to such a simulation as an induced future. It represents a hypothetical rollout of the learner’s predictive model under its current beliefs, abstracting away interaction with the external data stream. To distinguish this hypothetical evolution from real learning progression, we introduce a futures time index $s\in\mathbb{N}$ and denote steps within the simulated trajectory as superscripts on $X$, $Y$, and $Z$.
During the rollout, the learner generates targets from its own predictive distribution, while any components not represented by the learner are borrowed from the environment. Formally, given a history $H_{0:t}$ and state $Z_{t}$, the induced future evolves according to
$Y^{s}\sim q_{f}(\cdot\mid Z_{t}^{s-1},X^{s-1}),\quad X^{s}\sim q_{e}(\cdot\mid H_{0:s-1},Y^{s}),\quad Z_{t}^{s}\sim u^{\prime}(Z_{t}^{s-1},X^{s},Y^{s}),$ (3)
where $q_{e}$ is a hybrid distribution that treats the learner’s predictions as targets while borrowing components from the environment as needed, $u^{\prime}$ is the learner’s inference mode update function, and $s=(t+1,t+2,\dots)$. Progression along the futures index $s$ provides a principled means to analyse the learner independently of new observations. We emphasise that this futures distribution is entirely isolated from the interaction process.
We refer to an element of future histories as $H\in\mathcal{H}$, which is a sequence of future observation-output pairs $(X^{i},Y^{i})$ induced by the learner $H^{t+1:\infty}=((X^{t+1},Y^{t+1}),(X^{t+2},Y^{t+2}),\dots)$.
Definition 3.6 (Induced Future Distributions).
The induced future distribution of the learner at state $Z_{t}$ with realised history $H_{0:t}$ is the joint distribution
$q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$ (4)
defined by rolling out the inference-mode process in (3).
Thus, an induced future is a probability distribution over entire infinite sequences of future inputs and outputs, i.e., elements of $\mathcal{X}\times\mathcal{Y}$, which we denote by $(\mathcal{X}\times\mathcal{Y})^{\mathbb{N}}$. Because the learner’s state $Z_{t}$ and subsequent observations are random, the induced future distribution at each step $q(\cdot\mid Z_{t},H_{0:t})$ is a random variable taking values in $\mathcal{D}((\mathcal{X}\times\mathcal{Y})^{\mathbb{N}})$.

### v2

@@ Abstract
A fundamental challenge in developing general learning algorithms is their tendency to forget past knowledge when adapting to new data. Addressing this problem requires a principled understanding of forgetting; yet, despite decades of study, no unified definition has emerged that provides insights into the underlying dynamics of learning. We propose an algorithm- and task-agnostic theory that characterises forgetting as a lack of self-consistency in a learner’s predictive distribution over future experiences, manifesting as a loss of predictive information. Our theory naturally yields a general measure of an algorithm’s propensity to forget and shows that Bayesian learners are capable of adapting without forgetting. To validate the theory, we design a comprehensive set of experiments that span classification, regression, generative modelling, and reinforcement learning. We empirically demonstrate how forgetting is present across all deep learning settings and plays a significant role in determining learning efficiency. Together, these results establish a principled understanding of forgetting and lay the foundation for analysing and improving the information retention capabilities of general learning algorithms.
@@ 1 Introduction
Forgetting is a ubiquitous yet poorly understood phenomenon in machine learning (mccloskey1989catastrophic). When a learner updates its beliefs based on new observations, it often forgets prior knowledge. This leads to a degradation in performance on previous observations. Although this behaviour is well documented in continual learning (CL; kirkpatrick2017overcoming, kirkpatrick2017overcoming) and reinforcement learning (RL; atkinson2021pseudo, atkinson2021pseudo; khetarpal2022towards, khetarpal2022towards), forgetting also occurs in independent and identically distributed (i.i.d.) settings (chunking).
Most studies of forgetting come from the CL literature. Here, metrics usually track how performance degrades on earlier tasks after training on later tasks (chaudhry2018riemannian). Although widely adopted, these measures poorly capture forgetting and often conflate two distinct phenomena: backward transfer, where new learning improves performance on past tasks (benavides2022theory), and forgetting, where updates erode prior knowledge (jagielski2022measuring). This makes it challenging to distinguish between constructive and destructive adaptation.
In contrast, we study forgetting as a fundamental property of learning. It is a consequence of how any adaptive system updates its beliefs. CL, RL, and neural networks are not our focus. Instead, we treat them as instances of a broader phenomenon. We believe that before effective learning algorithms for CL can be developed, a precise conceptual understanding of forgetting must first be formulated. Therefore, our aim is to provide a general formalism that characterises forgetting in learning systems.
Our new conceptual foundation of forgetting is built on the following insight:
If a learner updates its predictions on data it already expects, that update cannot represent the acquisition of new information. Instead, it must represent the loss of previously acquired knowledge.
This allows us to give a precise and general definition of forgetting. We ask: What is forgetting? When and why does it occur? How does forgetting impact learning?
@@
    Formulation of learning over time: Inspired by hutter2005; dong2022simple; fortini2019quasi; fong2023martingale, we define a general theoretical formulation for reasoning about how learners acquire, retain, and lose capabilities during learning (§ 3).
@@
    Conceptualisation of forgetting: We formulate forgetting from a predictive perspective. Specifically, we define forgetting as a violation of self-consistency in the learner’s predictive distribution, providing a unified probabilistic foundation that generalises prior conceptions of forgetting (§ 4.2).
@@
    Operational measure: This conceptualisation naturally yields a measure of the propensity to forget, which we propose to assess the validity of our formalism (§ 4.2).
@@
    Empirical characterisation: We empirically study the propensity to forget in diverse environments and learning paradigms. These include regression, classification, generative modelling, CL, and RL. Our results confirm that forgetting dynamics conform to expected characteristics. They also reveal a trade-off between forgetting and training efficiency (§ 5).
@@ 2 Related work
Forgetting in CL.
Forgetting is often studied in the context of CL (de2021continual; wang2024comprehensive), often measured as the loss of performance on previous tasks (kirkpatrick2017overcoming; chaudhry2018efficient; chaudhry2018riemannian).
However, it is important to distinguish between two distinct effects: backwards transfer, where learning on new data improves performance on past tasks (benavides2022theory), and forgetting, where learning on new data degrades previously acquired knowledge (mccloskey1989catastrophic; jagielski2022measuring). Performance on previous tasks depends on both how current data informs the learner about past data (backwards transfer) and the extent to which updating on new tasks results in a loss of previous knowledge. Furthermore, most metrics are tailored to CL settings and do not easily accommodate arbitrary learning settings, domains, or learners. Thus, we suggest it is critical to develop a learner-agnostic definition of forgetting that can apply to CL and non-CL settings.
Various approaches have sought to quantify forgetting. In toy settings with known data distributions, forgetting can be measured relative to an oracle (lee2021continual). Others examine changes in internal representations (kim2025understanding) or model the trade-off between generalisation and forgetting (raghavan2021formalizing). Mechanistic accounts in associative memory models demonstrate how interference between correlated patterns leads to retrieval errors (hopfield_discrete; Amit). A related line of work studies models trained on their own generated outputs, typically in the context of generative model collapse (alemohammad2023self; shumailov2023curse; bertrand2023stability; scholten2025model). These studies implicitly touch on the phenomenon we formalise as forgetting: updating a model on self-generated targets can degrade previously encoded knowledge.
@@
Forgetting has also been studied in RL. Early work in continual RL noted the risk of knowledge degradation when agents adapt over long time horizons (ring1994continual; ring1997child). Recent surveys highlight that forgetting is a persistent challenge in RL (khetarpal2022towards). Empirical studies document several forms of forgetting, with forgetting measured in a manner similar to the CL counterpart, either by tracking the degradation of performance on previously learned tasks or by quantifying the distributional shift in the agent’s policy or value function over time (shenfeld2025rl). Value-based methods with function approximation often lose performance on earlier estimates (mnih2015human; van2018deep). This is closely related to the phenomenon of policy churn identified by (schaul2022phenomenon) in which the greedy policy of a value-based learner changes in a large portion of the input space after just a few updates. Policy gradient methods are also prone to overwriting earlier strategies during continual adaptation (ring1994continual; kirkpatrick2017overcoming). Replay buffers are widely used to mitigate these effects by reintroducing past experiences during training (random; fedus2020revisiting).
Characterisations of forgetting.
Across learning settings, forgetting is often characterised using a variety of different learner-specific mechanisms. For instance, model-centric views equate forgetting with parameter drift (mccloskey1989catastrophic; french1999catastrophic; rusu2016progressive; kirkpatrick2017overcoming; zenke2017continual; aljundi2018memory; masse2018alleviating; li2024spirf; zhao2023does), or policy drift (shenfeld2025rl), while accuracy-centric views characterise forgetting as performance decay on earlier tasks (kemker2018measuring; parisi2019continual; jagielski2022measuring; van2022three). Notably, these mechanism-specific views can be useful when these particular mechanisms are present within the class of learners of interest.
However, we again suggest that it is fruitful to conceive of forgetting from a general point of view that is agnostic to both the choice of learner, the loss, the domain, and the learning setting. For this reason, we adopt a general vantage and treat the predictive distribution as the object of interest, again embracing the predictive Bayes perspective discussed in § 3.2. Under this view, we develop a principled theory of forgetting based on predictive self-consistency that applies broadly across settings and learners. We further demonstrate the generality of our conceptualisation and show how it supports an empirical measure of forgetting that aligns with theoretical expectations.
@@ 3 Learning and inference processes
We present a general framework in which supervised learning, RL, and generative modelling are all specific cases of a single stochastic interaction process. The formalism draws inspiration from the agent-environment perspective from general RL (hutter2005; Lattimore2014; dong2022simple; Abel2023; hutter2024; kumar2025continual), but is adapted to emphasise conventions standard in machine learning (bishop2006pattern; goodfellow2016deep; fong2023martingale).
@@ 3.1 Preliminaries: Learning interaction
We formalise the data exchanged between a learner and the environment by defining an interface.
@@
Elements of $\mathcal{X}$ are observations made by the learner (e.g., numerical features, targets, etc.), and elements of $\mathcal{Y}$ are outputs emitted by the learner (e.g., actions, classifications, regression values, etc.).
@@
An element of $\mathcal{H}$ is a history $H\in\mathcal{H}$, which is a sequence $H_{0:t}=((X_{0},Y_{0}),\dots,(X_{t},Y_{t}))$ of $t+1$ observation-output pairs $(X_{i},Y_{i})$. For a sequence $A=(A_{0},A_{1},\dots,A_{t})$ and indices $0\leq i\leq j\leq t$, the subsequence between indices $i$ and $j$ is $A_{i:j}=(A_{i},A_{i+1},\dots,A_{j})$.
@@
A learner relative to interface $(\mathcal{X},\mathcal{Y})$ is a tuple $(\mathcal{Z},f,u,u^{\prime},p_{Z_{0}})$, where:
@@
Although a learner’s state could, in principle, be updated by a single function (as in work by dong2022simple, dong2022simple, abel2023convergence, abel2023convergence, and kumar2025continual, kumar2025continual), we distinguish between two update functions to capture different modes of learner evolution. During interaction with the environment, the learning-mode update $u$ governs the evolution of the entire state, including predictive parameters and auxiliary components, such as replay buffers. In contrast, the inference-mode update $u^{\prime}$ allows auxiliary components of the state, such as buffers or counters, to evolve while keeping the predictive parameters fixed. This distinction allows an observer to analyse the learner’s behaviour in both dynamic conditions, where beliefs are updated over time, and in static conditions, where beliefs remain fixed.
@@
The (learning) interaction process between an environment $(e,p_{X_{0}})$ and a learner $(\mathcal{Z},f,u,u^{\prime},p_{Z_{0}})$ relative to an interface $(\mathcal{X},\mathcal{Y})$ is defined by:
$\text{Initialisation }(t=0)\text{:}$ $Y_{0}=\bot,\quad X_{0}\sim p_{X_{0}},\quad Z_{0}\sim p_{Z_{0}},$
$\text{Interaction }(t\geq 1)\text{:}$ $Y_{t}\sim q_{f}(\cdot\mid Z_{t-1},X_{t-1}),$ (learner samples output)
$X_{t}\sim p_{e}(\cdot\mid H_{0:t-1},Y_{t}),$ (environment samples observation)
@@
This interaction process generates the stochastic processes $Y_{0}X_{0}Z_{0}Y_{1}X_{1}Z_{1}Y_{2}X_{2}Z_{2}\dots$
3.2 Predictive distributions
At any time during learning, the learner’s state encodes expectations about how future interactions will unfold. We can make these expectations explicit by asking the learner to internally simulate the future. Rather than waiting for new observations, we let the learner induce an infinite sequence of future interactions based on its current beliefs, using update $u^{\prime}$ throughout the induced process.
We refer to such a simulation as an induced future or predictive distribution. It represents a hypothetical rollout of the learner’s predictive model under its current beliefs. To distinguish this hypothetical evolution from real learning progression, we introduce a future time index $s\in\mathbb{N}$ and denote steps within the simulated trajectory as superscripts on $X$, $Y$, and $Z$.
During the rollout, the learner generates targets from its own predictive distribution. Formally, given a history $H_{0:t}$ and a state $Z_{t}$, the predictive distribution evolves according to
$Y^{s}\sim q_{f}(\cdot\mid Z_{t}^{s-1},X^{s-1}),\,X^{s}\sim q_{e}(\cdot\mid Y^{s},Z_{t}^{s-1}),\,Z_{t}^{s}\sim u^{\prime}(Z_{t}^{s-1},X^{s},Y^{s}),$ (3)
where $s\in\{t+1,t+2,\dots\}$, and $q_{e}$ is defined as the learner’s current distribution over observed data. Any component not modelled by the learner is instead sampled from the environment (e.g., inputs from a validation set). The futures distribution is entirely isolated from the interaction process.
We refer to an element of future histories $H\in\mathcal{H}$ as a sequence of future observation-output pairs $(X^{i},Y^{i})$ for $i\in s$, induced by the learner $H^{t+1:\infty}=((X^{t+1},Y^{t+1}),(X^{t+2},Y^{t+2}),\dots)$.
Definition 3.6 (Predictive Distributions).
The predictive distribution of the learner at state $Z_{t}$ with realised history $H_{0:t}$ is the joint distribution obtained by rolling out the inference-mode process in (3):
$q(H^{t+1:\infty}\mid Z_{t},H_{0:t}).$ (4)
Thus, an induced future is a probability distribution over entire infinite sequences of future inputs and outputs $(\mathcal{X}\times\mathcal{Y})^{\mathbb{N}}$. Because the learner’s state $Z_{t}$ and subsequent observations are random, the predictive distribution at each step $q(\cdot\mid Z_{t},H_{0:t})$ is a random variable taking values in $\mathcal{D}((\mathcal{X}\times\mathcal{Y})^{\mathbb{N}})$.